# chessformer.bib

### v1

@@ 5 Related Work
Independent work in the same vein as this paper by Ruoss et al. [ruoss2024grandmasterlevel] analyzes the effect of training transformer models at scale for chess, demonstrating that a transformer-based approximation of a strong oracle can be trained without domain-specific alterations to achieve grandmaster-level strength without search. Unlike this paper, they focus on distilling a search-based algorithm into a transformer model rather than optimizing for playing strength.
@@
Focusing instead on maximizing playing strength, Czech et al. [czech2023representation] design a lightweight transformer block for chess and show that the performance of neural chess models can be improved with carefully chosen input representations and value targets. Similar work in Go [wu2020accelerating] adds auxiliary training targets and enhanced input representations to improve convergence speed and final playing strength of the AlphaZero process.
@@
Other work interprets the inner workings of neural chess models through ablations, probing of internal states, and analysis of attention maps. McGrath et al. [mcgrath2021acquisition] analyze the acquisition of chess knowledge by AlphaZero during training, while Jenner et al. [lookahead] examine the role of attention maps and lookahead in an older transformer model trained by the Lc0 project.
6 Conclusion
@@
One key takeaway of this work is that, though the vanilla transformer is an effective generalist architecture, especially when trained at scale, there can be substantial benefit to exploring domain-specific enhancements and inductive biases. This holds true when resources are limited but especially in search-based applications where speed and accuracy are fungible. This idea is exemplified by work such as AlphaFold [alphafold], which predicts the structure of proteins by modeling them with a graph where edges correspond to relationships between residues.
7 Acknowledgements
This paper relied on extensive contributions from the Lc0 community. We are grateful to Twitch user KittenKaboodle for generously providing the hardware on which experiments were run. We are also thankful to David Elliott, Hunter Monroe, and Evan Engler for proofreading and providing insight. Any errors that remain are our own.
[…]
@@ Appendix A Model Architecture
A.1 Input Encoding
The input to our network is a sequence of 64 tokens of length 112 consisting of a concatenation of:
@@ A.1 Input Encoding
To generate token embeddings, we use a linear projection followed by adding and multiplying by learned offset vectors which are separate across tokens and depth. This gives the model absolute positional information and was found in initial experiments to incrementally improve model quality.
A.2 Body
The body of our models consists of a stack of encoder layers with Post-LN normalization and the initialization/gain scheme Deepnet [wang2022deepnet]. We use a fixed context length of 64, i.e., one token for each square on the board read from left to right then bottom to top, and the board is flipped with the side to move. We use Mish [misra2019mish] activations in the feedforward sublayer.
Following [palm], we omit biases in the QKV projections and omit centering and biases in the encoder normalization layers, finding this to increase training throughput without degrading quality. However, we retain biases in the feedforward and post-attention projections since removing them was found to degrade quality in initial experiments.
A.3 Output Heads
Similar to AlphaZero, our models have heads for policy and value predictions. However, we add several auxiliary policy and value targets to increase convergence speed, following research in Go by Wu [wu2020accelerating], [wu2024].
Policy Heads We introduce a new attention policy head based on scaled dot-product attention. Moves are encoded by the starting square of the piece moved and the destination square to which it moved. Given the sequence of tokens outputted by the body, we generate policy embeddings by applying a dense layer of depth equal to model’s embedding size. From this we generate via linear projection a set of query vectors corresponding to the starting square and a set of key vectors corresponding to the destination square, both with depth equal to the depth of the encoder body.
Logits for moves are calculated via standard scaled dot-product attention, described in Equation 1. We did not find improvement from more complicated self-attention variants, likely because the embedding size is large compared to the token count. To improve training stability, illegal moves are always masked.
@@ A.3 Output Heads
The “result” value head predicts the result of the game from among “win”, “draw”, and “loss” with cross entropy loss. The “q” and “short-term” value heads each predict three targets: a reward trained with L2 loss, a categorical distribution over rewards trained with cross-entropy loss, and the error between the predicted and true reward. Respectively, these latter two heads use the reward estimate produced during selfplay and an exponential moving average of future rewards with expected depth 6. The gradient of the reward prediction is detached when calculating the error loss to prevent the error head from affecting the reward prediction.
@@ Appendix B Train Losses
As described above, our models train on several auxiliary targets, which were found by the Katago team [wu2020accelerating], [wu2024] to accelerate model convergence and improve final performance. Here we describe the losses for these targets. The final loss function is a weighted sum of these terms.
@@
    where $\pi$ is the policy target and $\hat{\pi}$ is the model’s prediction of $\pi$ and $c_{\text{pol}}=1$.
@@
  • L2 value: predicts a scalar reward produced by selfplay.
@@
Appendix C Playstyle
We notice that our largest model, CF-240M, has a very humanlike understanding of the game, which allows it to assess positions in ways that elude top minimax-based engines. One such position arose in a game between Lc0 and Stockfish at the 23rd Computer Chess Championship Rapid tournament (see here), where the former used a slightly older transformer model with 190 million parameters. Lc0 managed to construct a fortress from a lost position, leading the game into a draw. We show a position from this game in Figure 3(a). CF-240M assigns the only drawing move, Kg7, a policy of 92.8%, and evaluates the position as drawn (2.2%W, 76.3%D, 21.5%L). Stockfish assigned the same position a +7 evaluation at 11 billion nodes, and its evaluation stayed above +3 for over 100 moves.
Another position demonstrating this humanlike positional understanding arose in the 21st Computer Chess Championship Rapid tournament (see here), where Lc0 used a medium-sized transformer model of around 100 million parameters. Lc0 played b4, trapping the white rook on b3. CF-240M assigns this move the highest policy at 37% and evaluates the position as slightly winning (40.3%W, 54.1%D, 5.6%L).
If the pawn on b2 is removed and this trapping idea is no longer feasible, the evaluation becomes much more drawish (17.1%W, 77.7%D, 5.2%L), despite black gaining a material advantage. Additionally, the policy on b4 is reduced to 4.76%, and Rc7 becomes the top policy move. Lc0 ended up winning the game, with Stockfish’s evaluation only going negative several moves later despite searching hundreds of millions of nodes.
Our models’ positional understanding seems to extend to long-term planning. As shown in Figure 3(c), CF-240M picks the right move at each of white’s turns in a famous king walk played by GM Nigel Short as white against GM Jan Timman. The line continues 31. Kh2 Rc8 32. Kg3 Rce8 33. Kf4 Bc8 34. Kg5.
(a) Fortress detection: CF-240M finds the only drawing move, Kg7, and evaluates the position as drawn.
(b) Trapped piece detection: when white has a pawn on b2, CF-240M opts to trap the white rook on b3 by pushing b4. However, if the pawn on b2 is not present, it abandons this idea.
(c) Long-term planning: CF-240M finds a famous king walk that was played out by Nigel Short.
Figure 3: Humanlike play: our transformer models exhibit a humanlike understanding of the game, detecting positional ideas that elude top minimax-based engines.
Appendix D Other Techniques Tried
In addition to the techniques reported here, we tried several architectural modifications that did not bear fruit. For example, we found replacing the feed-forward layer with a Gated Linear Unit [glu] to slightly degrade performance when model FLOPS were held constant. Using a sparsely gated [moe] or soft [softmoe] mixture of experts in place of the feed-forward layer also did not improve performance.

### v2

@@ VII Related Work
Independent work in the same vein as this paper by Ruoss et al. [7] trains transformers at scale for chess, demonstrating that a transformer-based approximation of a strong oracle can achieve grandmaster-level strength without search. Unlike this paper, they focus on distilling a search-based algorithm into a transformer model rather than optimizing for playing strength.
@@
Czech et al. [9] design a lightweight transformer block for chess and show that the performance of neural chess models can be improved with carefully chosen input representations and value targets. Similar work in Go [10], [11] adds auxiliary training targets and enhanced input representations to accelerate learning in the AlphaZero process.
@@
Other work interprets the inner workings of neural chess models through ablations, probing of internal states, and analysis of attention maps. McGrath et al. [12] analyze the acquisition of chess knowledge by AlphaZero during training, while Jenner et al. [13] examine the role of attention maps and lookahead in an older iteration of transformer models trained by the authors.
VIII Conclusion
@@
Though the vanilla transformer is an effective generalist architecture, as we have shown, there can be substantial benefit to exploring domain-specific enhancements and inductive biases like effective position representations. This holds true in resource-limited scenarios but especially in search-based applications like computer chess where speed and accuracy are interchangeable. This idea is exemplified by work such as AlphaFold [14], which predicts the structure of proteins by modeling them with a graph where edges correspond to relationships between residues.
The fungibility of quantity and quality has featured prominently in recent work aiming to improve the reasoning capabilities of language models by encouraging them to delineate their thought process before giving a final answer. For example, chain-of-thought prompting [15] prompts a language model with example reasoning paths, increasing accuracy at the expense of additional computational cost. OpenAI’s o1 model [16] takes this idea further and is trained to perform a large reasoning step before giving its final answer. In the domain of automated theorem-proving, AlphaProof [17] solves olympiad-level math problems by searching through potential proofs with a model specialized for theorem proving.
We hope our work, which demonstrates the potential usefulness of domain-specific enhancements and shows that in some cases deep learning can succeed where search fails, can provide lessons for this new search-based reasoning paradigm.
@@ Appendix A Model Architecture
A-A Input Encoding
The input to our network is a sequence of 64 tokens: one token for each square on the board read from left to right then bottom to top. The board is flipped with the side to move. Each input token has length 112 and consists of a concatenation of:
@@ A-A Input Encoding
To generate token embeddings, we apply a linear projection to the input tokens, then add and multiply by learned offset vectors which are separate across tokens and depth. This gives the model absolute positional information and was found in initial experiments to incrementally improve model quality.
A-B Body
The body of our models consists of a stack of encoder layers with Post-LN normalization and the initialization/gain scheme DeepNorm [18]. We use a fixed context length of 64. We use Mish [19] activations in the feedforward sublayer, and all projections in the output heads are followed by this activation unless otherwise stated.
Following [20], we omit biases in the QKV projections and omit centering and biases in the encoder normalization layers, finding this to increase training throughput by around 10% without degrading model quality. However, we retain biases in the feedforward and post-attention projections since removing them degraded quality in initial experiments.
A-C Output Heads
Similar to AlphaZero, our models have heads for policy and value predictions. However, we add several auxiliary policy and value targets to increase convergence speed, following research in Go [10], [11].
Policy Heads We introduce a new policy head based on a modified self-attention mechanism. Moves are encoded by the starting square and destination square of the piece moved. Given the sequence of tokens outputted by the body, we generate policy embeddings by applying a dense layer of depth equal to the model’s embedding size. From this we generate, via linear projection, a set of query vectors corresponding to the starting square and a set of key vectors corresponding to the destination square, both with depth equal to the depth of the encoder body.
Logits for moves are calculated via scaled dot product as described in Eq. 1. The result is a 64x64 matrix representing all possible traversals from one square on the chessboard to another. These traversals are sufficient to represent all moves that can occur within the rules of chess, with the exception of promotions. When a pawn advances to the last rank of the board, it must be promoted to a knight, bishop, rook, or queen. To represent these special moves, we apply a linear projection to the key vectors representing the promotion rank, generating an additive bias for each possible promotion piece. This bias is then applied to the logits representing all possible traversals between the penultimate rank and the promotion rank to generate additional logits for each possible promotion.
To generate the final policy vector, we apply a softmax over all logits, masking illegal moves to increase training stability. We did not see performance gain from using any of the position representations described in Section II, likely because the embedding size is much larger than the token count.
@@ A-C Output Heads
The “result” value head predicts the result of the game from among win, draw, and loss with cross-entropy loss. The “q” and “short-term” value heads each predict three targets: a reward trained with L2 loss, a categorical distribution over rewards trained with cross-entropy loss, and the error between the predicted and true reward. Respectively, these latter two heads use the reward estimate produced during self-play and an exponential moving average of future rewards with expected depth 6. The gradient of the reward prediction is detached when calculating the error loss to prevent the error head from affecting the reward prediction.
@@ Appendix B Train Losses
As noted above, our models train on several auxiliary targets, which were found to accelerate model convergence and improve final performance in Go [10], [11]. Here we describe the losses for the main and auxiliary targets. The final loss function is a weighted sum of these terms.
@@
    where $\pi$ is the policy target, $\hat{\pi}$ is the model’s prediction of $\pi$, and $c_{\text{pol}}=1$.
@@
  • L2 value: predicts a scalar reward produced by self-play.
@@
Appendix C Other Techniques Tried
In addition to the techniques reported here, we tried several architectural modifications that did not bear fruit. For example, replacing the feed-forward layer with a Gated Linear Unit [21] slightly degraded quality at constant model FLOPS. Using a sparsely gated [22] or soft [23] mixture of experts in place of the feed-forward layer also did not improve performance.
Acknowledgment
This paper relied on extensive contributions from the Lc0 community. We are grateful to Twitch user KittenKaboodle for generously providing the hardware on which experiments were run, and to Github user jkormu for implementing the attention visualization. We are also thankful to David Elliott, Hunter Monroe, Denys Ivanov, and Evan Engler for proofreading and providing insight. Any errors that remain are our own.
Training code and templates for Fig. 1, Fig. 2, and Fig. 5 were generated with help from Github Copilot, and much of the text was checked for spelling and writing quality with GPT-4.